Equal weights of methane and hydrogen are mixed in an empty container at 25°C. the fraction of the total pressure exerted by hydrogen is
a. 1/2
b. 8/9
c. 1/9
d. 16/17

Grade:11

2 Answers

Kevin Nash
9 years ago
Sol. Let x grams of each hydrogen and methane are mixed, Moles of H2 = x/2 Moles of CH4 = x/16 ⇒ Mole-fraction of H2 = (x/2)/(x/2 + x/16) = 8/9 ⇒ (Partial pressure of H_2)/(Total pressure) = mole-fraction of H2 = 8/9
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Let x grams of each hydrogen and methane are mixed,
Moles of H2 = x/2
Moles of CH4 = x/16
⇒ Mole-fraction of H2 = (x/2)/(x/2 + x/16)
= 8/9
⇒ (Partial pressure of H_2)/(Total pressure)
= mole-fraction of H2
= 8/9

Thanks and Regards